OpenStudy (anonymous):

245% of 24.2 is what number?

OpenStudy (jonnyvonny):

Mmk, you do this the same way as you would to fine any percentage. part/total *100 = %

OpenStudy (jonnyvonny):

(part/total)*100 = percentage

OpenStudy (anonymous):

I got a long decimal number

OpenStudy (jonnyvonny):

Mmk lets break it down

OpenStudy (anonymous):

Well I did 245/24.2 and got 10.12 along with other numbers

OpenStudy (jonnyvonny):

First, we want to know what number, when compared to 24.2, times 100, gives 245

OpenStudy (jonnyvonny):

245%*

OpenStudy (jonnyvonny):

So part x total 24.2 *100 = 245

OpenStudy (jonnyvonny):

\[\frac{ x }{ 24.2 }*100 = 245%\]

OpenStudy (jonnyvonny):

Just solve for "x".

OpenStudy (anonymous):

From multiplying 24.2 by 100 I got 2420

OpenStudy (jonnyvonny):

Are you with me, or is that confusing?

OpenStudy (jonnyvonny):

No, to get rid of *100, you /100

OpenStudy (anonymous):

Hmm it is kind of confusing, but I'm still trying

OpenStudy (texaschic101):

or... turn your percent into a decimal....245% = 2.45 " of " means multiply so.. 2.45(24.2) = your answer

OpenStudy (anonymous):

Oh that makes more sense

OpenStudy (anonymous):

How will I know when to divide and when to multiply?

OpenStudy (texaschic101):

well...in math word problems, " of " means multiply there is also another way to do percent problems by using the percent formula percent formula : is/of = %/100 now look at your problem.. 245% of 24.2 is what number.. % = 245 " of " = 24.2 " is " = a number we will call x now just sub x/24.2 = 245/100 -- since this is a proportion, we now cross multiply 100x = (24.2)(245) 100x = 5929 -- divide by 100 x = 5929/100 x = 59.29 <== your answer